# Reconsidering the Performance of GAE in Link Prediction

### v3

@@ 5.3. Ablation Study
Table 5. Ablation study for Optimized-GAE, evaluating key design choices covering: (1) Input Representation; (2) Convolution Layers; (3) Layer Design (impact of residual connections and MPNN encoder linearity).
@@ 5.3.1. Input Nodes Representations
Explanation: Table 4 clearly supports this. A higher $I_{S/F}$ index typically indicates that LE are more beneficial. For instance, on the ogbl-ppa dataset, which lacks informative raw features ($I_{S/F}\approx 1.000$), using LE substantially improves performance from approximately $21.37$ to $78.41$. Conversely, for a feature-rich dataset like Citeseer ($I_{S/F}\approx 0.292$), using its strong RF alone gives a result of $92.40$, which is better than the $87.21$ achieved if LE are used. While the overall success of LE also depends on graph size and density (see Observation 2), the $I_{S/F}$ index provides an important first assessment. It helps determine if the node features are sufficient or if the model needs to focus more on the graph structure.
@@
Explanation: The major impact of initialization is clear from Table 5. For instance, while Optimized-GAE uses LEs with an orthogonal initialization and achieves a Hits@20 score of $94.43$ on ogbl-ddi, performance drops sharply with less structured initializations. Specifically, using an “all-ones initialization,” where initial embedding vectors are all ones, results in a score of only $2.13$. Similarly, using “random initialization,” where embeddings are initialized with values selected uniformly from [1,1][-1,1] (a method likely to break any initial orthogonality), greatly reduces performance to $13.84$. This large drop in performance highlights the importance of orthogonal initialization. Notably, even after training, these learnable embeddings often remain close to orthogonal. Specifically, on ogbl-ddi, the average absolute cosine similarity was 0.03 (std. dev. 0.04) before training and 0.07 (std. dev. 0.07) after training. This reflects the important role of common neighbor information in link prediction, as discussed in Section 4.1.
@@
Explanation: Table 5 also shows the benefits of learnability. While using ’fixed orthogonal embeddings’ (initialized orthogonally but not updated during training) on ogbl-ddi achieves a Hits@20 score of $77.61$, this is much lower than the $94.43$ achieved by the full Optimized-GAE with learnable embeddings. This performance gap highlights that allowing the embeddings to be fine-tuned enables the model to use more than just common neighbor information and capture detailed representations specific to the graph structure.
@@ 5.3.2. Key Architectural Modules of Graph Autoencoders
Observation 5: While GCN serve as a effective default for the Optimized-GAE, alternative convolutional architectures exhibit varying performance trade-offs, underscoring the importance of layer selection tailored to dataset characteristics.
@@
Explanation: The significant advantage of using linear encoder layers is supported by the “Nonlinear MPNN Layers” ablation presented in Table 5 (section “Layer Design”). When non-linear activation functions are added back between the MPNN layers in the Optimized-GAE’s encoder, performance consistently and often substantially drops on multiple datasets. For example, on ogbl-ddi, the result falls from $94.43$ to $66.77$. These results strongly support our design choice of using linear propagation in the encoder. This allows the model to effectively capture and keep pairwise relational signals and structural information, as discussed in Section 4.1.
@@ 5.3.3. Hyperparameter Tuning: Network Depth and Hidden Dimension
Table 6. Comparison of time complexity per training batch and total runtime (seconds) on evaluation sets across various models and datasets. Here, $B$ denotes batch size, and $h$ the hashing sketch size for BUDDY and MPLP+. Detailed complexity analysis is provided in Appendix B.
@@
MPNN Layers $\{2,\,3,\,4\}$ 4 4 4 2 4 2 3
Hidden Dimension $\{256,\,512,\,1024\}$ 1024 1024 512 1024 512 512 256
Batch Size [2048, 65536][2048,\,65536] 2048 4096 4096 8192 16384 65536 65536
Dropout $\{0.2,0.4,0.6\}$ 0.6 0.6 0.4 0.6 0.2 0.2 0.2
MLP Layers $\{2,4,5,8\}$ 4 2 2 8 5 5 5
@@ 6. Efficiency
In addition to its strong performance, Optimized-GAE is also highly scalable due to its simple architecture. Models like GAE do not require explicit calculation of pairwise information, which gives them an inherent efficiency advantage over models that use complex methods to include pairwise information. For example, SEAL requires subgraph extraction and DRNL calculation for each node (Zhang and Chen, 2018), while BUDDY precomputes subgraph-related information and structural features for each batch (Chamberlain et al., 2023), causing significant overhead. Table 6 compares the time complexity per training batch and the actual runtime per training epoch on the same machine, highlighting that our GAE-based approach offers scalability benefits over models with complex pairwise methods, even when using a large hidden dimension and deeper layers.
@@ 7. CONCLUSION
This work demonstrates that a holistically optimized GAE can achieve SOTA link prediction performance, rivaling sophisticated contemporary models often with superior efficiency. Our success, achieved through systematic and validated enhancements to GAE’s input representation strategies, core architectural designs, and meticulous hyperparameter tuning, underscores the critical importance of re-evaluating simple baseline models. This study not only provides a robustly tuned baseline and practical design guidelines but more broadly highlights that advancements in graph learning can significantly benefit from the thorough understanding and refinement of well-established architectures.
Appendix A Experimental Details
Datasets To evaluate the performance of Optimized-GAE, we leverage a variety of widely used benchmark datasets. These datasets are sourced from two established repositories: Planetoid (Sen et al., 2008) and OGB (Hu et al., 2020). From the Planetoid collection, we utilize Cora, Citeseer, and Pubmed, while the OGB repository provides ogbl-collab, ogbl-ddi, ogbl-ppa, and ogbl-citation2. Together, these datasets cover a range of graph structures and relationship patterns, enabling a comprehensive evaluation of model performance.
Computing infrastructure Our model is built with DGL (Wang et al., 2019) and PyTorch (Paszke et al., 2019). Experiments are conducted on Nvidia 4090 GPUs, except for the ogbl-citation2 dataset, which is run on Nvidia A800 GPUs due to its higher computational requirements.
Training process We set the upper bound for training epochs to 500, slightly higher than previous settings, as we observed that performance continues to improve on certain datasets up to this limit. This extended training duration is feasible due to the relatively high training speed of GAE. A 1:3 negative sampling ratio was employed (three negative edges per positive edge). All results are computed from runs conducted with 5 different random seeds.
Model Hyperparameters We tune the hyperparameters using a bayesian optimization, selecting values that yield the best validation performance. Table 7 summarizes the hyperparameter settings.
Appendix B Complexity analysis
First, we analyze the time complexity of Optimized-GAE. A typical MPNN layer such as a GCN involves a round of message passing and a nonlinear transformation. This results in a total time complexity of $O(|E|d+|V|d^{2})$. However, Optimized-GAE skips the nonlinear transformation, reducing the complexity to $O(|E|d)$. Additionally, training involves running an MLP on a batch of edges with a complexity of $O(Bld^{2})$. For space complexity, MPNN layers have $O(|E|+|V|d)$, where the second term corresponds to node embeddings. MLP predictors have $O(Bd+ld^{2})$, with the first term for edge embeddings and the second for the weight matrix.
We further analyze the time complexity of recent models in Table 6. For each link, SEAL extracts a subgraph and runs an MPNN, leading to a batch complexity of $O(B|E|d+B|V|d^{2})$. BUDDY only runs a lightweight MPNN on the full graph once instead and computes hashing sketches, resulting in a complexity of $O(|E|d+|E|h)$, followed by an MLP predictor with complexity $O(Bd^{2}+Bh)$. MPLP+ propagates orthogonal sketches in addition to the standard GAE model, introducing an extra complexity of $O(|E|h)$.
@@ Acknowledgments
We thank the anonymous reviewers of CIKM 2025 for their time and insightful feedback, which helped improve this paper.

### v4

@@ 5.1.2. Implementation Details
Table 6. Ablation study for Optimized GAE, evaluating key design choices covering: (1) Input Representation; (2) Convolution Layers; (3) Layer Design (impact of residual connections and MPNN encoder linearity).
@@
5.2. Main Results
The main results are presented in Table 3. Our model achieves SOTA on 3 datasets and top-2 on 4 datasets. The consistently strong performance reflects an average improvement of 91.5% over the original GAE baseline. For instance, on the structure-reliant ogbl-ppa and ogbl-ddi datasets, where our input representation technique is particularly effective, performance improves from 18.67% to 78.41% and from 37.07% to 94.43%. Furthermore, Optimized GAE shows an average improvement of 6.4% over NCN (Wang et al., 2024b), a competitive SOTA model, surpassing it on six out of seven datasets. Notably, on ogbl-ppa, our model’s Hits@100 score of 78.41% not only exceeds the strong MPLP+ baseline by a significant 20.2% margin but also outperforms the OGB leaderboard leader by 2.4%. These results support our main claim: a properly optimized GAE can match or exceed the performance of current models that use more complex methods, validating GAE’s potential in current link prediction tasks.
5.3. Ablation Study
@@ 5.3.1. Input Nodes Representations
Explanation: Table 5 clearly supports this. A higher $I_{S/F}$ index typically indicates that LEs are more beneficial. For instance, on the ogbl-ppa dataset, which lacks informative raw features ($I_{S/F}\approx 1.000$), using LE substantially improves performance from approximately $21.37$ to $78.41$. Conversely, for a feature-rich dataset like Citeseer ($I_{S/F}\approx 0.292$), using its strong RF alone gives a result of $92.40$, which is better than the $87.21$ achieved if LEs are used. While the overall success of LE also depends on graph size and density (see Observation 2), the $I_{S/F}$ index provides an important first assessment. It helps determine if the node features are sufficient or if the model needs to focus more on the graph structure.
@@
Explanation: The major impact of initialization is clear from Table 6. For instance, while Optimized GAE uses LEs with an orthogonal initialization and achieves a Hits@20 score of $94.43$ on ogbl-ddi, performance drops sharply with less structured initializations. Specifically, using an “all-ones initialization,” where initial embedding vectors are all ones, results in a score of only $2.13$. Similarly, using “random initialization,” where embeddings are initialized with values selected uniformly from [1,1][-1,1] (a method likely to break any initial orthogonality), greatly reduces performance to $13.84$. This large drop in performance highlights the importance of orthogonal initialization. Notably, even after training, these learnable embeddings often remain close to orthogonal. Specifically, on ogbl-ddi, the average absolute cosine similarity was 0.03 (std. dev. 0.04) before training and 0.07 (std. dev. 0.07) after training. This reflects the important role of common neighbor information in link prediction, as discussed in Section 4.1.
@@
Explanation: Table 6 also shows the benefits of learnability. While using ’fixed orthogonal embeddings’ (initialized orthogonally but not updated during training) on ogbl-ddi achieves a Hits@20 score of $77.61$, this is much lower than the $94.43$ achieved by the full Optimized GAE with learnable embeddings. This performance gap highlights that allowing the embeddings to be fine-tuned enables the model to use more than just common neighbor information and capture detailed representations specific to the graph structure.
@@ 5.3.2. Key Architectural Modules of Graph Autoencoders
Observation 5: While GCN serve as a effective default for the Optimized GAE, alternative convolutional architectures exhibit varying performance trade-offs, underscoring the importance of layer selection tailored to dataset characteristics.
@@
Explanation: The significant advantage of using linear encoder layers is supported by the “Nonlinear MPNN Layers” ablation presented in Table 6 (section “Layer Design”). When non-linear activation functions are added back between the MPNN layers in the Optimized GAE’s encoder, performance consistently and often substantially drops on multiple datasets. For example, on ogbl-ddi, the result falls from $94.43$ to $66.77$. These results strongly support our design choice of using linear propagation in the encoder. This allows the model to effectively capture and keep pairwise relational signals and structural information, as discussed in Section 4.1.
@@ 5.3.3. Hyperparameter Tuning: Network Depth and Hidden Dimension
Table 7. Comparison of time complexity per training batch and total runtime (seconds) on evaluation sets across various models and datasets. Here, $B$ denotes batch size, and $h$ the hashing sketch size for BUDDY and MPLP+.
@@ 6. Efficiency
In addition to its strong performance, Optimized GAE is also highly scalable due to its simple architecture. Models like GAE do not require explicit calculation of pairwise information, which gives them an inherent efficiency advantage over models that rely on more complex mechanisms.
The theoretical basis for this efficiency is clear from the model’s time complexity. The Optimized GAE encoder uses linear layers, resulting in a complexity of just $O(|E|d)$, while the MLP decoder operates on a batch of $B$ edges with complexity $O(Bld^{2})$. This is significantly more efficient than competing methods. For instance, SEAL requires extracting and processing a unique subgraph for each link, leading to a high batch complexity of $O(B|E|d+B|V|d^{2})$. BUDDY must compute hashing sketches over the graph, giving it a complexity of $O(|E|d+|E|h)$ followed by a predictor of $O(Bd^{2}+Bh)$, while MPLP+ adds an extra $O(|E|h)$ for propagating orthogonal sketches on top of its GAE-style backbone. Table 7 compares the theoretical time complexity per training batch and the actual runtime on our evaluation sets. These results highlight that our GAE-based approach offers substantial scalability benefits over models with complex pairwise methods, even when using a large hidden dimension and deeper layers.
@@ 7. CONCLUSION
This work demonstrates that a holistically optimized GAE can achieve SOTA link prediction performance, rivaling sophisticated contemporary models often with superior efficiency. Our success, achieved through systematic and validated enhancements to GAE’s input representation strategies, core architectural designs, and meticulous hyperparameter tuning, underscores the critical importance of re-evaluating simple baseline models. This study not only provides a robustly tuned baseline and practical design guidelines but more broadly highlights that advancements in graph learning can significantly benefit from the thorough understanding and refinement of well-established architectures. Looking ahead, the power and efficiency of our Optimized GAE makes it a strong candidate for future extension to more complex domains such as dynamic and heterogeneous graphs.
@@ Acknowledgments
This work is supported by the National Key R&D Program of China (2022ZD0160300) and National Natural Science Foundation of China (62276003).